The owners of a marina bought 6

canoes to be used as rentals at a cost of ​$400

per canoe. They borrowed 80
​%
of the total cost for 10

months at 7 and one half
​%
interest. Find the amount due using simple interest.

Answer by stanbon(75887)

The owners of a marina bought 6 canoes to be used as rentals at a cost of ​$400per canoe. They borrowed 80​% of the total cost for 10months at 7 and one half % interest. Find the amount due using simple interest.
--------
Amt. borrowed:: 0.8*6*400 = 0.8*2400 = $1920
-------
Amt Due = 1920(1+rt) = 1920+1920*0.075*(10/12) = $2040
------------------
Cheers,
Stan H.
------------------